OpenStudy (anonymous):

How do you solve cos(pi/24) using Half-Angle formulas, and leaving in simplified form?

OpenStudy (mani_jha):

\[\cos (\theta/2)=(\sqrt{1+\cos (\theta)})/2\] Using this, first find cos(pi/12) by taking theta=pi/6 then find cos(pi/24) by taking theta =pi/12 ok?

OpenStudy (anonymous):